### v1

# Geodesic structure of spacetime near singularities

Abstract
Geodesic flows emanating from an arbitrary point $\mathscr{P}$ in a manifold $\mathscr{M}$ carry important information about the geometric properties of $\mathscr{M}$. These flows are characterized by Synge’s world function and van Vleck determinant - important bi-scalars that also characterize quantum description of physical systems in $\mathscr{M}$. If $\mathscr{P}$ is a regular point, these bi-scalars have well known expansions around their flat space expressions, quantifying local flatness and equivalence principle. We show that, if $\mathscr{P}$ is a singular point, the scaling behavior of these bi-scalars changes drastically, capturing the non-trivial structure of geodesic flows near singularities. This yields remarkable insights into classical structure of spacetime singularities and provides useful tool to study their quantum structure.
I Introduction
An important first step in understanding the role of quantum effects in regions of strong gravitational fields is to understand the behavior of geodesic flows - behavior of geodesics emanating from a single point - in vicinity of the singularity. This is so because many important features of classical and quantum dynamics depend on this behavior; see, for instance, Gutzwiller (1990) for an extensive textbook discussion on this point. While the structure of these flows is well understood in geodesically convex neighborhoods of regular spacetime points, not much is known about their behavior in regions containing curvature singularities. This is to be expected since covariant Taylor expansions that work well in the former case breaks down in the latter, and hence the analysis of geodesic flows directly in terms of curvature related quantities can not be done for generic singularities. We derive here, for the first time, explicit behavior of Synge’s world function $\Omega(x,y)$ and van Vleck determinant $\Delta(x,y)$ - two important bi-scalars characterizing geodesic flows - in the vicinity of FLRW and Kasner like singularities. In the process, we improve upon an old result by Buchdahl Buchdahl and Warner (1980) who calculated the series expansion of $\sqrt{|\Omega|}$” in FLRW and obtained a series whose terms can be shown to diverge as one approaches the singularity at $t\to 0$. This obviously un-physical result, as we shall see, is an artifact of expanding a wrong function (and manifests, in fact, even in the Minkowski limit!), and was presumably not noticed by Buchdahl. Here, we focus instead on $\Omega(x,y)$ itself, which we take to be the only meaningful quantity.
The results exhibit a remarkable and non-trivial modification in behavior of these bi-scalars near the singularities. We elaborate on these results and put them in the broader context of quantum gravity. Over the past decade, several research works on quantum spacetime have recognized the relevance of Synge’s world function and van Vleck determinant in reconstructing and “effective” spacetime metric $q_{ab}(x;x_{0})$ Kothawala (2013),Kothawala (2023a),Kothawala (2023b), Padmanabhan (1997), Pesci (2025). The effective metric description already yields some very interesting insights into local properties of spacetime in a geodesically convex neighborhood of a regular event $\mathscr{P}_{0}$. Our results here enable extending these insights when $\mathscr{P}_{0}$ represents a curvature singularity. These results will be presented in a separate work Kothawala which, in fact, provided an important motivation for the current work.
To keep our discussion streamlined and focus clear, we give definitions of the key quantities and summarize our key results at the end of this section (see below). Rest of the paper will provide a detailed exposition on various aspects of the results, and is organized as follows: Section II gives the key derivation leading to systematic representations of $\Omega(x,y)$ and $\Delta(x,y)$ in FLRW and Bianchi spacetimes. In Section III, we use these results to discuss the geodesic and causal structure of these spacetimes, focussing on the limit where the base point is taken as the singularity as $T\to 0$. Finally, we conclude with a brief discussion in Section IV.
I.1 Synge’s world function and van Vleck determinant
Synge’s world function and van Vleck determinant are two important bi-scalars that characterize measurements in curved spacetimes, and also appear ubiquitously in quantum field theory. Two-point (and higher) correlators of quantum fields in curved spacetime have a universal expansion in the world function near the coincidence limit. These expansions form the backbone of the point-splitting program for regularization in curved spacetime, and has been used developed and studied extensively by de Witt, Cristensen, Parker and others in deriving covariant expansions of effective lagrangians in curved spacetime, covariant formulation of path integral etc. The use of the world function as a fundamental object to study spacetime dates back to Synge, where it is defined operationally as half the square of the geodesic distance between the two spacetime points (denoted by $\Omega(x,x^{\prime})$). The spacetime metric $g_{ab}(x)$ can be constructed as the coincidence limit of $\nabla_{a}\nabla_{b}\Omega(x,y)$, i.e.
$g_{ab}=\lim_{x^{\prime}\longrightarrow x}\nabla_{a}\nabla_{b}\Omega(x,x^{\prime})=[\nabla_{a}\nabla_{b}\Omega(x,x^{\prime})]$ (1)
Similarly, higher order coincidences gives us information about the spacetime curvature. Thus world function can be used to study tensor expansions on a given spacetime manifold in a covariant manner. The only hindrance to it’s use (apart from existence of multiple geodesics connecting two points) is the non-existence of any closed-form expression for geodesic interval as a function of spacetime coordinates even in highly symmetric spacetimes. In fact, to the best of our knowledge, closed form expressions are known only in maximally symmetric and plane-wave spacetimes Hari and Kothawala (2021),Harte and Drivas (2012). It would surprise the in-cognoscenti that the simplest and most widely encountered spacetimes such as Schwarzschild and FLRW have no closed form expression for $\Omega$. In this work, we derive a systematic representation for the world function in FLRW and Bianchi spacetimes, in a form that is applicable near singularities in these spacetimes.
Figure 1: The most basic non-local observables in an arbitrary curved manifold are Synge’s world function and the van Vleck determinant. The former characterizes single geodesics (left), while the latter measures the density of geodesics emanating from a given point (right). $\lambda$ is the affine parameter along the geodesic. See text for more details.
van Vleck determinant is a bivector of central importance that can be computed using the world function directly. It appears at various places, like fluctuation factor in semi-classical approximation to path integrals in curved spacetime, reconstruction of spacetime while defining q-metric using van Vleck determinant, heat kernel expansions in curved spacetime, etc. Given a geodesic with endpoints $x$ and $z$, It is defined as,
$\Delta(x,z):=-\frac{1}{\sqrt{-g(x)}}\det\left(\frac{\partial^{2}\Omega(x,z)}{\partial x^{a}\partial z^{b}}\right)\frac{1}{\sqrt{-g(z)}}$ (2)
Where $\Omega(x,z)$ is Synge’s world function. More intuitively, it’s a quantitative measure of density of geodesics emanating from a given point as shown in fig.1.
I.2 Summary of results
The table 1 gives the summary of the coincidence and singularity limits obtained for the various biscalars in FLRW and Bianchi Type I spacetime. We obtained the closed-form expression for Synge’s world function for FLRW and Kasner (Schwarzschild) spacetime in the form of an infinite series. The limits of $\Delta$ and $\Box\Delta^{1/2}$ and $K=\textrm{Tr}(K_{ab})$ are computed to study the geodesic and causal structure near the curvature singularity. The lightcones structure is illustrated in fig. 3 and fig. 4 by plotting null geodesics along with $\Omega=$ constant surfaces about different points in the neighbourhood of FLRW and Kasner (Schwarzschild) singularity.
Bitensor FLRW Bianchi Type I (Schwarzschild)
Coincidence Singularity Coincidence Singularity
$\Delta(t,T)$ $\textrm{matter}:1+\frac{\varepsilon^{2}}{9T^{2}}+\cdots$ $\textrm{radiation}:1+\frac{\varepsilon^{2}}{8T^{2}}+\cdots$ $\textrm{matter}:\frac{t}{27T}$ $\textrm{radiation}:\left(\frac{t}{T}\right)^{3/2}\frac{1}{(\log(t/T))^{3}}$ $1+O(\varepsilon^{4})$ $\frac{5t^{1/3}}{27T^{1/3}}+\frac{10}{27}+\cdots$
$\Box\Delta^{1/2}(t,T)$ $K(t,T)$ $\textrm{matter}:\frac{2}{9T^{2}}$ $\textrm{radiation}:\ \ 0$ $\frac{3}{\varepsilon}+\frac{\ddot{a}}{a}\varepsilon$ $\textrm{matter}:\frac{2267}{2187\sqrt{21}t^{2}}$ $\textrm{radiation}:\ \ \ 0$ $\frac{3q}{t}$ $\frac{3}{\varepsilon}-\frac{8\varepsilon^{3}}{1215T^{4}}+$... $\frac{8}{3t}$
$\Omega(t,T)$ $-\frac{(t-T)^{2}}{2}+\sum_{j=1}^{\infty}\frac{\ell^{2j}}{(2j)!}\left[\frac{d^{2j-1}}{d\alpha^{2j-1}}\Bigl\{\frac{d\Omega(\alpha)}{d\alpha}(G)^{-2j}\Bigr\}\right]_{\alpha=0}$ $\sum_{i,j,k=0}^{\infty}\mathcal{C}_{ijk}(x-X)^{i}(y-Y)^{j}(z-Z)^{k}$
Table 1: Summary of important limits in FLRW and Bianchi Type I spacetimes
II Representations of $\Omega$ and $\Delta$ in FLRW and Bianchi spacetimes
While the definition of the World function $\Omega(x,y)$ is straightforward, and it is numerically equal to one half of the squared geodesic interval between $x$ and $y$, it is extremely difficult to obtain a closed form expression for it even in the simplest of spacetimes. The difficult lies in determining the functional dependence of the $\Omega(x,y)$ on the end-point coordinates $x$ and $y$. It is due to this reason that one often has to resort to coincidence limit expansions of the World function in specific coordinates. However, such an expansion might be rendered useless simply due to choice of bad coordinates, or near spacetime singularities, which is the case of interest for us in this work. We show that the World function in the vicinity of a class of singularities (FLRW or Kasner type) can be represented in an expansion that is mathematically well behaved.
II.1 FLRW spacetime
In this section we consider the case of FLRW spacetime and compute the world function for a timelike geodesic. The FLRW metric with an arbitrary scale factor $a(t)$ is given by,
$ds^{2}=-dt^{2}+a^{2}(t)(dx^{2}+dy^{2}+dz^{2})$ (3)
As mentioned in the Introduction, an attempt to calculate the series expansion of $\sqrt{\Omega}$ by solving the Hamilton-Jacobi equation, first made by Buchdahl Buchdahl and Warner (1980), leads to spurious and unphysical divergences in the limit $t\to 0$. The only meaningful quantity to calculate is the Synge world function $\Omega$ itself, which, for a timelike geodesic, is given by, $\Omega=-\tau^{2}/2$, where $\tau$ is the proper time along the geodesic. A general approach to compute proper time as a function of end points amounts to solving the geodesic equation ($\nabla_{\bm{u}}\bm{u}=0$) for a timelike geodesic, which is in general a difficult task. However, owing to the presence of spacelike Killing vector fields in the FLRW spacetime, the computation of proper time can be reduced to a quadrature, yielding the following expressions Roberts (1993),
$\tau(t,T;\alpha)$ $=$ $\int_{T}^{t}\frac{a(t^{\prime})}{\sqrt{\alpha^{2}+a^{2}(t^{\prime})}}dt^{\prime}$ (4)
$\ell(t,T;\alpha)$ $=$ ${\alpha}\int_{T}^{t}\frac{dt^{\prime}}{a(t^{\prime})\sqrt{\alpha^{2}+a^{2}(t^{\prime})}}=\alpha G(\alpha)$ (5)
where $\tau(t,T;\alpha)$ is the proper time as a function of coordinate time of end points and integration constant $\alpha$, $\ell(t,T;\alpha)$ is the geodesic distance on the three-space of constant curvature. Now $\Omega$ can be rewritten using ($d\tau^{2}/dt^{\prime}=2\tau(d\tau/dt^{\prime})$) and We eliminate the integration constant $\alpha$ for $\ell$ using Lagrange-Good formula Abdesselam (2003) in single variable to get $\tau(t,T;\ell)$ and arrive at the following expansion for $\Omega(t,T;\ell)$ (see Appendix IV for details),
$\Omega(t,T;\ell)$ $:=$ $\Omega(0)+\sum_{j=1}^{\infty}\frac{\ell^{2j}}{(2j)!}\left[\frac{d^{2j-1}}{d\alpha^{2j-1}}\left[\frac{d\Omega}{d\alpha}(G)^{-2j}\right]\right]_{\alpha=0}$ (6)
where $\Omega(0)=-(t-T)^{2}/2$, This expansion in $\ell$ can be written in a closed form by identifying the following expansions,
$\cosh(\ell D_{\alpha}[(G)^{-1}\Omega(\alpha)])-1$ $:=$ $\sum_{j=1}^{\infty}\frac{(\ell D_{\alpha})^{2j}[(G)^{-2j}\Omega(\alpha)]_{\alpha=0}}{(2j)!}$ (7)
$\sinh(\ell D_{\alpha}[(G)^{-1}(G^{-1})^{\prime}\Omega(\alpha)])$ $:=$ $\sum_{j=1}^{\infty}\frac{(\ell D_{\alpha})^{2j-1}[(G)^{-(2j-1)}(G^{-1})^{\prime}\Omega(\alpha)]_{\alpha=0}}{(2j-1)!}$ (8)
where $D_{\alpha}=d/d\alpha$, so that the final expression for the world function can be written concisely as follows.
$\Omega(t,T;\ell)=\Omega(0)+\cosh(\ell D_{\alpha}[G^{-1}\Omega(\alpha)])+\ell\sinh(\ell D_{\alpha}[G^{-1}(G^{-1})^{\prime}\Omega(\alpha)])$ (9)
The (9) is a non-perturbative closed-form expression in terms of the geodesic length $\ell$ in geodesic distance on the three-space of constant curvature. The complication of further simplification arises because of the time-dependent scale factor $a(t)$, which is kept open for generality. As another self-consistency check, it can be verified that (IV) satisfies the HJ equation up to order $\mathcal{O}(\ell^{4})$ with;
$\Omega(t,T;\ell)=-\frac{(t-T)^{2}}{2}+\frac{(t-T)}{2I_{2}}\ell^{2}+\left[\frac{(t-T)I_{4}-\left(I_{2}\right)^{2}}{8\left(I_{2}\right)^{4}}\right]\ell^{4}+\mathcal{O}(\ell^{6})$ (10)
It is worth emphasizing the generality of this result, since $\ell$ appearing in eq.(9) is the geodesic distance on the three-space of constant curvature. Thus the expression for $\Omega(t,T;\ell)$ holds for flat, open and closed slicing of FLRW spacetime.
An immediate biscalar that we can construct from the $\Omega(t,T;\ell)$ is the van Vleck determinant $\Delta(t,T)$ which characterizes the geodesic spread about a given point, If we consider the expression (10) derived earlier, then the van Vleck determinant for FLRW spacetime up to $\mathcal{O}(\ell^{2})$ is given by,
$\Delta(t,T)\simeq\frac{1}{a(t)^{3}a(T)^{3}F(t,T)^{3}}+\frac{20F(t,T)^{3}Q(t,T)-F^{(1,1)}(t,T)}{2a(t)^{3}a(T)^{3}F(t,T)^{5}}\ell^{2}+\mathcal{O}(\ell^{4})$ (11)
with
$F(t,T)$ $=$ $\frac{2}{t-T}I_{2}(a)$ (12)
$Q(t,T)$ $=$ $\left[\frac{(t-T)I_{4}(a)-I_{2}(a)^{2}}{8I_{2}(a)^{4}}\right]$ (14)
where $I_{n}(a_{i})$ is defined by (31). We now consider the limiting behaviour of $\Delta(t,T)$ along a timelike geodesic in coincidence and singularity limit in matter and radiation dominated FLRW spacetime. The expansion in $\epsilon=T-t<<T$ (coincidence limit) and (Singularity limit) $T<<t,T\to 0$ are obtained as follows,
  1. 1.
    Matter-dominated FLRW spacetime
    $\textrm{Coincidence}:\ \ \ \ \Delta(t,T)$ $\simeq$ $1+\frac{\epsilon^{2}}{9T^{2}}+\mathcal{O}(\epsilon^{3})$ (15)
    $\textrm{Singularity}:\ \ \ \ \Delta(t,T)$ $\simeq$ $\frac{t}{27T}+$... (16)
  2. 2.
    Radiation-dominated FLRW spacetime
    $\textrm{Coincidence}:\ \ \ \ \Delta(t,T)$ $\simeq$ $1+\frac{\epsilon^{2}}{8T^{2}}+\mathcal{O}(\epsilon^{3})$ (17)
    $\textrm{Singularity}:\ \ \ \ \Delta(t,T)$ $\simeq$ $\left(\frac{t}{T}\right)^{3/2}\frac{1}{(\log(t/T))^{3}}+$... (18)
The different divergent terms appearing in singularity limits are akin to what happens near caustics Harte and Drivas (2012). It can be independently verified that coincidence limit expansion for matter and radiation dominated FLRW spacetime agrees with, $\Delta\sim 1+R_{ab}\Omega^{a}\Omega^{b}/6$ behavior, whereas the singularity limit expansion depends on the specific choice of $a(t)$. The higher order derivatives of $\Delta(t,T)$ such as $\Box\Delta^{1/2},\Box\Box\Delta^{1/2}$... contribute crucially to Seeley-deWitt coefficients in a heat kernel expansion and are often expressed in terms of geometric invariants in coincidence limits, such as Álvarez and Anero (2022),
$[\Box\Delta^{1/2}]$ $=$ $R/6$ (19)
$\left[\Box\Box\Delta^{1/2}\right]$ $=$ $\frac{1}{5}\Box R+\frac{1}{36}R^{2}-\frac{1}{30}R_{ab}^{2}-\frac{1}{30}R_{abcd}^{2}$ (20)
The singularity and coincidence limits of $\Box\Delta^{1/2}$ turns out to be computationally tractable in our case and are given by,
  1. 1.
    Matter-dominated FLRW spacetime
    $\textrm{Coincidence}:\ \ \ \Box\Delta^{1/2}(t,T)$ $=$ $\frac{2}{9T^{2}}$ (21)
    $\textrm{Singularity}:\ \ \ \Box\Delta^{1/2}(t,T)$ $\simeq$ $\frac{2267}{2187\sqrt{21}t^{2}}$ (22)
  2. 2.
    Radiation-dominated FLRW spacetime
    $\textrm{Coincidence \& Singularity}:\ \ \ \Box\Delta^{1/2}(t,T)$ $=$ $0$ (23)
The coincidence limit of $\Box\Delta^{1/2}$ diverges as we approach the singularity, whereas the singularity limit ($T\to 0$) remains finite. For radiation-dominated FLRW spacetime ($R=0$), $\Box\Delta^{1/2}$ vanishes for both coincidence and singularity limits!
II.2 Bianchi Type I singularities
In this section, we consider the world function for the Bianchi type I metric, which is an exact vacuum solution to Einstein’s equations representing a spatially homogeneous (independent of $x,y,z$) but an anisotropic universe with different rates of expansion/contraction in different spatial directions. Bianchi type I solutions have been used extensively in studying cosmological singularities, BKL oscillations, and gravitational chaos near singularities Misner (1969)Belinsky et al. (1970)Belinsky et al. (1982). A Bianchi type I metric is defined as follows,
$ds^{2}=-dt^{2}+a^{2}_{1}(t)dx^{2}+a_{2}^{2}(t)dy^{2}+a_{3}^{2}(t)dz^{2}$ (24)
where the $a_{1},a_{2},a_{3}$ are scale factors providing a staightforward anisotropic generalization to FLRW spacetime. It turns out that this anisotropy can be handled analytically using three spacelike killing vector corresponding to translations in $(x,y,z)$, giving us three integration constants to deal with. The world function for this case is defined as follows,
$\Omega(t,T;\alpha_{1},\alpha_{2},\alpha_{3})=-\frac{\tau^{2}}{2}=-\frac{1}{2}\left[\int_{T}^{t}\left(1-\sum_{i=1}^{3}\frac{\alpha_{i}^{2}}{a_{i}^{2}(t)+\alpha_{i}^{2}}\right)^{1/2}dt\right]^{2}$ (25)
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are integration constants. The eqn. (25) can be inverted using Lagrange-Good formula Abdesselam (2003) for multivariables to give a series expansion in (xX),(yY),(zZ)(x-X),(y-Y),(z-Z), where
$(x_{i}-X_{i})=\int_{T}^{t}\frac{\alpha_{i}}{a_{1}(t^{\prime})(a_{1}^{2}(t^{\prime})+\alpha_{i}^{2})^{1/2}}dt^{\prime}=\alpha_{i}\ \phi_{i}(\alpha_{i})$ (26)
to give,
$\Omega$ $=$ $\sum_{i,j,k=0}^{\infty}\mathcal{C}_{ijk}(x-X)^{i}(y-Y)^{j}(z-Z)^{k}$ (27)
$\textrm{where}\ \ \ \mathcal{C}_{ijk}$ $=$ $\frac{1}{i!j!k!}\left[\frac{\partial^{i+j+k}}{\partial\alpha_{1}^{i}\partial\alpha_{2}^{j}\partial\alpha_{3}^{k}}\left(\Omega(\alpha_{1},\alpha_{2},\alpha_{3})\phi_{1}^{-i}\phi_{2}^{-j}\phi_{3}^{-k}\det\left(\delta_{ab}+\alpha_{b}\frac{\partial\log\phi_{a}}{\partial\alpha_{b}}\right)\right)\right]_{\alpha_{1},\alpha_{2},\alpha_{3}=0}$ (28)
where $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $a,b=1,2,3$. The $\mathcal{O}((x_{i}-X_{i})^{4})$ expansion for $\Omega$ using (27) gives,
$\Omega$ $=$ $-\frac{(t-T)^{2}}{2}+\frac{(t-T)}{2}\sum_{i=1}^{3}I_{2}(a_{i})(x_{i}-X_{i})^{2}+\sum_{i,j=1,i\neq j}^{3}\ \frac{(t-T)I_{2}(a_{i}a_{j})-I_{2}(a_{i})I_{2}(a_{j})}{(2I_{2}(a_{i})I_{2}(a_{j}))^{2}}(x_{i}-X_{i})^{2}(x_{j}-X_{j})^{2}$ (29)
$+$ $\ \ \sum_{i=1}^{3}\frac{(t-T)I_{4}(a_{i})-(I_{2}(a_{i}))^{2}}{8(I_{2}(a_{i}))^{4}}(x_{i}-X_{i})^{4}+\mathcal{O}\left((x_{i}-X_{i})^{6}\right)$ (30)
where,
$I_{n}(a_{i}):=\int_{T}^{t}\frac{dt^{\prime}}{a(t^{\prime})^{n}}$ (31)
A potentially interesting case for theoretical interest is the Kasner limit of the Schwarzschild metric, which corresponds to the following choice of scale factors,
$a_{1}(t)$ $=$ $(t/t_{0})^{2/3}$ (32)
$a_{2}(t)$ $=$ $(t/t_{0})^{2/3}$ (33)
$a_{3}(t)$ $=$ $(t/t_{0})^{-1/3}$ (34)
The world function to $\mathcal{O}\left((x_{i}-X_{i})^{2}\right)$ for this case is given by,
$\Omega_{\textrm{schw}}(t,T;\ell_{\perp},z-Z)$ $=$ $-\frac{(t-T)^{2}}{2}+\frac{(t-T)}{6t_{0}((t_{0}/T)^{1/3}-(t_{0}/t)^{1/3})}\ell^{2}_{\perp}$ (35)
$+$ $\frac{5(t-T)}{6(t(t/t_{0})^{2/3}-T(T/t_{0})^{2/3})}(z-Z)^{2}+\mathcal{O}\left((x_{i}-X_{i})^{4}\right)$ (36)
Where $\ell^{2}_{\perp}=(x-X)^{2}+(y-Y)^{2}$. It is easy to verify that eqn(35) satisfies the necessary coincidence limits of a world function i.e. $\Omega_{\rm schw}]=[\nabla_{a}\Omega_{\rm schw}]=0\ ,\ [\nabla_{a}\nabla_{b}\Omega_{\rm schw}]=g_{ab}=\text{diag}(-1,(t/t_{0})^{2/3},(t/t_{0})^{2/3},(t/t_{0})^{-1/3})$ and gives a self consistent series expansion of $\Omega$ in Schwarzschild case. Furthermore, we calculated the van Vleck determinant for Kasner limit of Schwarzschild spacetime, which shows different scaling behaviour in singular and coincidence limits.
$\Delta_{\text{schw}}$ $\simeq$ $1+\mathcal{O}(\varepsilon^{4})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ :\ \textrm{Coincidence limit}$ (37)
$\Delta_{\text{schw}}$ $\simeq$ $\frac{5t^{1/3}}{27T^{1/3}}+\frac{10}{27}+\mathcal{O}(T^{1/3})\ \ \ \ :\ \ \textrm{Singularity limit}$ (38)
The first non-trivial contribution in coincidence limit expansion of $\Delta_{\textrm{schw}}$ comes at order $\mathcal{O}(\varepsilon^{4})$ because order $\mathcal{O}(\varepsilon^{2})$ is identically zero for Schwarzschild spacetime ($R_{ab}=0$). The singularity limit of $\Delta_{\textrm{schw}}$ shows a $1/T^{1/3}$ divergence as compared to $1/T$ divergence we observed in the case of matter-dominated FLRW spacetime; the different scaling behaviour can be attributed to the presence of non-zero shear and rotation in Kasner (Schwarzschild) spacetime as compared to an isotropic expansion in matter-dominated FLRW spacetime.
III Geodesic and causal structure near singularity
Figure 2: A local region of matter-dominated FLRW spacetime is shown with t=t=constant and $\Omega=$ constant anchored at a point $P(T,X,Y,Z)$. The extrinsic curvature $K_{ab}$ of $\Omega$=constant hypersurfaces characterizes the local expansion of these surfaces about point $P$. The t=t=constant hypersurfaces follow Hubble’s law, whereas $\Omega=$constant hypersurfaces obey a generalized version of Hubble’s law Kothawala (2021).
The equi-geodesic surfaces are defined as surfaces of constant bi-scalar $\Omega(p,\mathscr{P})$ about a fiducial spacetime point. These surfaces constitute a fundamental geometric structure which has been extensively explored in Stargen and Kothawala (2015)Kothawala (2013). Let us now consider the global behaviour of these equi-geodesic surface $\Sigma(\kappa)$ defined by $\{p\ |\ \Omega(p,P)=-\kappa\ ,\ p\in\mathcal{M}\}$ in a FLRW spacetime with one of the points of world function $\Omega$ anchored at $P=(T,X,Y,Z)$. The equations for $\Sigma(\kappa)$ and lightcones are given by,
$\Omega$ $=$ $-\kappa=-\frac{(t-T)^{2}}{2}+\frac{(t-T)}{2I_{2}}\ell^{2}$ (39)
$\ell$ $=$ $\pm\int_{T}^{t}\frac{dt^{\prime}}{a(t^{\prime})}$ (40)
We plot the set of equations (39) in the vicinity of the singularity $\mathscr{P}$ about different points $P$ translated in time to study the causal structure near the singularity. Since FLRW spacetime is homogeneous, we consider $X,Y,Z=0$ without loss of generality. The lightcones flatten out in the FLRW case (see fig 3) but still remain asymptotes to the time-like hypersurfaces. We find a contrasting behavior for Schwarzschild geometry, the right figure in 3 shows light cones drawn at different points $P$ along the radial infalling geodesic of a point particle in the vicinity of the Schwarzschild singularity. The light cones begin to shrink and converge as we approach the Schwarzschild singularity along the contracting dimension. The Schwarzschild metric in Kasner form is given by,
$ds^{2}\simeq-d\tau^{2}_{\circ}+\tau_{\circ}^{-2/3}dz^{2}+\tau_{\circ}^{4/3}d\mathbb{S}_{2}$ (41)
Where $\tau_{\circ}$ is the proper time of a radial infalling observer and $z=t$ is the Schwarzschild time, which becomes spacelike inside the black hole.
The equi-geodesic surfaces can be locally characterized by their extrinsic curvature $K$ as demonstrated by one of the author in Stargen and Kothawala (2015). We now use the $\mathcal{O}(\ell^{2})$ expression for the world function of FLRW metric to derive an explicit expression for the extrinsic curvature $K_{ab}$ for $\Omega=\text{constant}$ surfaces in this spacetime. The normal $n_{a}$ to the $\Omega=-\kappa=\textrm{constant}$ surface is defined as,
$n_{a}=\frac{\nabla_{a}\Omega}{2\sqrt{-\Omega}}$ (42)
The extrinsic curvature $K_{ab}$ and trace $K$ for this normal evaluates to,
$K_{ab}=\nabla_{a}n_{b}=\left(\frac{1}{(t-T)F(t,T)}+a(t)\dot{a}(t)\right)\begin{pmatrix}0&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}$ (43)
$K=\textrm{Tr}(K_{ab})=\frac{3}{(t-T)F(t,T)a^{2}(t)}+\frac{3\dot{a}(t)}{a(t)}$ (44)
The trace $K=K(t,T)$ as defined above is a biscalar. The coincidence limit expansion of eq.(44) in expansion parameter $\varepsilon=t-T$ gives,
$\textrm{Coincidence}:\ \ \ \ K=\frac{3}{\varepsilon}+\frac{\ddot{a}(t)}{a(t)}\varepsilon+\mathcal{O}(\varepsilon^{2})$ (45)
The $\varepsilon^{-1}$ divergence term is a universal divergence that is independent of the scale factor $a(t)$ and can be attributed to the flat space limit. The second term in eq.(45) corresponds to the tidal trace and matches with another result obtained by one of the author in a more general setting Stargen and Kothawala (2015). The universal divergent term in $K$ can be renormalized by subtracting out $K_{0}=-3/\varepsilon$ to give,
$\tilde{K}=K-K_{0}\simeq\frac{\ddot{a}(t)}{a(t)}\varepsilon+\mathcal{O}(\varepsilon^{2})$ (46)
Thus, if we have a local construction of $\Omega=\textrm{constant}$ surface about a point $(T,X,Y,Z)$ in its convex neighbourhood, the extrinsic curvature of that surface depends on the second derivative of the scale factor! It is quite interesting to note that this is distinct from the Hubble expansion term $3\dot{a}/a$.
On the other hand, the singularity limit, $T\to 0,t\to 0^{+}$, $K$ evaluates to
$\textrm{Singularity}:\ \ \ \ K\simeq\frac{3q}{t}$ (47)
The leading term appearing in this expression is nothing but the expansion scalar $\Theta$ for $t=$ constant surfaces defined by $\Theta:=d(\ln V)/dt={3q}/{t}$ with $V\propto a^{3}\propto(t/t_{0})^{3q}$. This peculiar result can be understood from the fact that, in FLRW spacetimes, the $\Omega=\textrm{constant}$ surfaces and $t=\textrm{constant}$ surfaces tend to merge smoothly as we approach the singularity; this is illustrated by explicit plots of these surfaces for matter-dominated case given in Fig.3.
The extrinsic curvature and its trace for Kasner (Schwarzschild) evaluates to,
$K_{ab}=\nabla_{a}n_{b}$ $=$ $\sum_{a,b=1}^{3}\left(\frac{1}{I_{2}(a_{i})}+a_{i}(t)\dot{a}_{i}(t)\right)\delta_{ab}$ (48)
$\textrm{Singularity}:K$ $=$ $\frac{8}{3t}$ (49)
$\textrm{Coincidence}:K$ $=$ $\frac{3}{\varepsilon}-\frac{8\varepsilon^{3}}{1215T^{4}}+.$... (50)
where $a_{i}$ are defined as in eq.(32). The behaviour of lightcones in this case should be contrasted with the Kasner case for Schwarzschild singularity (see 4 and right half of Fig. 3) where lightcones close up along the contracting dimension (here $z$), reducing the region of spacetime that remains in causal contact for the timelike observer as singularity is approached.
Figure 3: The equi-geodesic surfaces and lightcones are constructed using biscalar $\Omega(x^{a},X^{a})$ about different points $P$ in the spacetime. Left: Matter-dominated FLRW spacetime with $p=2/3$, Right: Kasner form of Schwarzschild singularity. The zig-zag line shows the curvature singularity in both cases.
Figure 4: The lightcones near a Kasner (Schwarzschild) singularity stretches due to non-zero shear and anisotropic nature of spacetime, whereas lightcones in FLRW spacetime expands isotropically. Left: Lightcones in Kasner (Schwarzschild) elongate (left) along the expanding direction and squeeze (right) along the contracting direction (here $z$). Right: General comparison of lightcone structure between Kasner (Schwarzschild) and FLRW (matter-dominated) spacetime.
IV Discussion
In this work, we have derived a series representation for Synge’s world function for FLRW and Bianchi type I spacetime. Interestingly, the key structural properties of the world function are captured through integrals of $1/a(t)^{2n}$, $n\in\mathbb{N}$, between the end-points of a given timelike geodesic. Such dependence was also observed by Buchdahl long back Buchdahl and Warner (1980) in expansion of $\sqrt{\Omega}$, but this expansion turns out to be bad near singularities due to spurious, unphysical divergences. Our representation, on the other hand, gives well-defined limits if one of the points is singular.
As an important concomitant, we have computed the singular and coincidence limits of the van Vleck determinant for FLRW (matter-dominated and radiation) and Bianchi class of spacetimes, focussing in particular on the Kasner limit of the Schwarzshild near singularity. The van Vleck determinant shows a divergent power law behaviour in the singular limit for matter-dominated FLRW and Kasner cases, whereas a logarithmic divergence for the radiation-dominated FLRW case. We provide some further insights into the geodesic structure close to singularities by investigating the behavior of extrinsic curvature for the equi-geodesic ($\Omega=$ constant) surfaces anchored at a point close to the singularity. The trace of extrinsic curvature gives us the expansion of $\Omega=$ constant surfaces, and we compare this with the expansion of t=t=constant surface; the differences are interesting and shown in Fig. 2. The contrasting behaviour of coincidence and singularity limits of biscalars suggests that the spacetime neighbourhood near a singularity is qualitatively different from a regular region of spacetime. This can have implications for small scale structure of quantum spacetime. Several recent works have demonstrated the important role of bi-tensors in characterizing small scale structure of spacetime; Kothawala (2013),Kothawala (2023a),Kothawala (2023b). It is a major consensus that the Planck length $\ell_{p}$ acts as a universal regulator at small scales and gives rise to non-trivial relics Kothawala (2023b).
Appendix A Calculational details in FLRW spacetime
we can reduce the solution of calculating proper time to quadrature Roberts (1993) and get the following expressions,
$\tau(t,T;\alpha)$ $=$ $\int_{T}^{t}\frac{a(t^{\prime})}{\sqrt{\alpha^{2}+a^{2}(t^{\prime})}}dt^{\prime}$ (51)
$\ell(t,T;\alpha)$ $=$ ${\alpha}\int_{T}^{t}\frac{dt^{\prime}}{a(t^{\prime})\sqrt{\alpha^{2}+a^{2}(t^{\prime})}}=\alpha G(\alpha)$ (52)
where $\alpha$ is an integration cosntant, Now $\Omega$ can be rewritten as (using $d\tau^{2}/dt^{\prime}=2\tau(d\tau/dt^{\prime})$),
$\Omega(t,T;\alpha)$ $=$ $-\frac{\tau^{2}}{2}=-\int_{T}^{t}\left(\tau\frac{d\tau}{dt^{\prime}}\right)dt^{\prime}$ (53)
$=$ $-\int_{T}^{t}\left[\int_{T}^{t^{\prime}}\frac{a(t_{1})}{\sqrt{a^{2}(t_{1})+\alpha^{2}}}dt_{1}\right]\frac{a(t^{\prime})}{\sqrt{\alpha^{2}+a^{2}(t^{\prime})}}dt^{\prime}$ (54)
So now we have $\Omega(\alpha)$ to be expanded using the Lagrange-Good formula Abdesselam (2003) in powers of $\ell$. Thus expansion in $\ell$ given as,
$\Omega(t,T;\ell):=\Omega(0)+\sum_{k=1}^{\infty}\frac{\ell^{k}}{k!}\left[\frac{d^{k-1}}{d\alpha^{k-1}}({\Omega}(\alpha)^{\prime}G^{-k}(\alpha))\right]_{\alpha=0}$ (55)
where,
$\frac{d\Omega}{d\alpha}=\alpha\int_{T}^{t}\int_{T}^{t^{\prime}}\frac{dt_{1}dt^{\prime}a(t_{1})a(t^{\prime})}{\sqrt{\alpha^{2}+a(t_{1})^{2}}\sqrt{\alpha^{2}+a(t^{\prime})^{2}}}\left[\frac{1}{\alpha^{2}+a(t^{\prime})^{2}}+\frac{1}{\alpha^{2}+a(t_{1})^{2}}\right]=\alpha F(\alpha)$ (56)
It is easy to confirm that $\Omega(0)=-(t-T)^{2}/2$, which makes sense in $\ell\longrightarrow 0$ limit. Since functions $\Omega,G,F$ are all even functions of $\alpha$, only even terms contribute in the expansion.
$\Omega(t,T;\ell)=-\frac{(t-T)^{2}}{2}+\sum_{j=1}^{\infty}\frac{\ell^{2j}}{(2j)!}\left[\frac{d^{2j-1}}{d\alpha^{2j-1}}(\alpha F)(G)^{-2j}\right]_{\alpha=0}$ (57)
where $F$ and $G$ are defined in eqns (56) and (51). Let us now consider $j=1$ term explicitly, The expansion for $G$ is,
[…]
where,
$I_{n}(a)=\int_{T}^{t}\frac{dt_{1}}{a(t_{1})^{n}}$ (59)
Let us now consider expansion for $F(\alpha)$,
$I=\int_{T}^{t}\int_{T}^{t^{\prime}}\frac{dt_{1}dt^{\prime}a(t_{1})a(t^{\prime})}{(\alpha^{2}+a(t_{1})^{2})^{1/2}(\alpha^{2}+a(t^{\prime})^{2})^{3/2}}=\int_{T}^{t}\left[\int_{T}^{t^{\prime}}\frac{a(t_{1})dt_{1}}{(\alpha^{2}+a^{2}(t_{1}))^{1/2}}\right]\frac{dt^{\prime}a(t^{\prime})}{(\alpha^{2}+a^{2}(t^{\prime}))^{3/2}}dt^{\prime}$ (60)
Now, the integral in the square bracket can be expanded in a series as,
[…]
Thus,
[…]
Therefore expression for $F$ is given by,
[…]
where $(t_{1}\leftrightarrow t^{\prime})$ is given by,
$(t_{1}\leftrightarrow t^{\prime})$ $=$ $\int_{T}^{t}\left[\int_{T}^{t^{\prime}}\frac{a(t_{1})dt_{1}}{(\alpha^{2}+a^{2}(t_{1}))^{3/2}}\right]\frac{dt^{\prime}a(t^{\prime})}{(\alpha^{2}+a^{2}(t^{\prime}))^{1/2}}dt^{\prime}$ (64)
[…]
Therefore, the complete expression for $F$ is given by,
[…]
Let’s calculate lowest order contribution in $\alpha$ to $(-\alpha F(G^{-1})^{2})$, i.e. $n=k=m=0$ in series for $G$ and $F$.
Therefore,
$G_{0}$ $=$ $I_{2}(t)=\int_{T}^{t}\frac{dt_{1}}{a^{2}(t_{1})}$ (70)
$F_{0}$ $=$ $F|_{m,k=0}=\int_{T}^{t}\left[\frac{I_{0}(t^{\prime})}{a^{2}(t^{\prime})}+I_{2}(t^{\prime})\right]dt^{\prime}$ (71)
$=$ $\int_{T}^{t}\left[\frac{t^{\prime}-T}{a^{2}(t^{\prime})}+\int_{T}^{t^{\prime}}\frac{dt_{1}}{a^{2}(t_{1})}\right]dt^{\prime}$ (72)
$=$ $(t^{\prime}-T)|_{T}^{t}\int_{T}^{t}\frac{dt^{\prime}}{a^{2}(t^{\prime})}-\int_{T}^{t}\int_{T}^{t^{\prime}}\frac{dt_{1}}{a^{2}(t_{1})}dt^{\prime}+\int_{T}^{t}\int_{T}^{t^{\prime}}\frac{dt_{1}}{a^{2}(t_{1})}dt^{\prime}$ (73)
$=$ $(t-T)\int_{T}^{t}\frac{dt^{\prime}}{a^{2}(t^{\prime})}dt^{\prime}$ (74)
Therefore $\mathcal{O}(\ell^{2})$ contribution to $\Omega$ is given by,
$\Omega(t,T;\ell)=-\frac{(t-T)^{2}}{2}+\frac{(t-T)}{2}\left[\int_{T}^{t}\frac{dt^{\prime}}{a^{2}(t^{\prime})}\right]^{-1}\ell^{2}+\mathcal{O}(\ell^{4})$ (75)
Consider the expression,
$\Omega(t,T;\ell)$ $=$ $\Omega(0)+\sum_{j=1}^{\infty}\frac{\ell^{2j}}{(2j)!}\left[\frac{d^{2j-1}}{d\alpha^{2j-1}}\Bigl\{(\alpha F)(G^{-1})^{2j}\Bigr\}\right]_{\alpha=0}$ (76)
$=$ $\Omega(0)+\sum_{j=1}^{\infty}\frac{\ell^{2j}}{(2j)!}\left[\frac{d^{2j-1}}{d\alpha^{2j-1}}\Bigl\{\frac{d\Omega}{d\alpha}(G^{-1})^{2j}\Bigr\}\right]_{\alpha=0}$ (77)
Consider the series separately.
$\frac{d^{2j-1}}{d\alpha^{2j-1}}\left(\frac{d\Omega}{d\alpha}(G^{-1})^{2j}\right)$ $=$ $\frac{d^{2j-1}}{d\alpha^{2j-1}}\left(\frac{d}{d\alpha}(\Omega(G^{-1})^{2j})-\Omega\ 2j(G^{-1})^{2j-1}(G^{-1})^{\prime}\right)$ (78)
$=$ $\frac{d^{2j-1}}{d\alpha^{2j-1}}\left(\frac{d}{d\alpha}(\Omega(G^{-1})^{2j})+2j\ \Omega(G^{-1})^{2j-1}(G^{-1})^{\prime}\right)$ (79)
$=$ $\frac{d^{2j}}{d\alpha^{2j}}(\Omega(G^{-1})^{2j})+2j\frac{d^{2j-1}}{d\alpha^{2j-1}}[\Omega(G^{-1})^{2j-1}(G^{-1})^{\prime}]$ (80)
Thus,
$\Omega(t,T;\ell)-\Omega(0)=\sum_{j=1}^{\infty}\frac{\ell^{2j}}{(2j)!}\frac{d^{2j}}{d\alpha^{2j}}(\Omega(\alpha)(G)^{-2j})+\ell\sum_{j=1}^{\infty}\frac{\ell^{2j-1}}{(2j-1)!}\frac{d^{2j-1}}{d\alpha^{2j-1}}[\Omega(\alpha)(G)^{-2j+1}((G)^{-1})^{\prime}]$ (81)
Note that we have the following identification,
$\cosh(\ell D_{\alpha}[(G)^{-1}...])-1$ $:=$ $\sum_{j=1}^{\infty}\frac{(\ell D_{\alpha})^{2j}[(G)^{-2j}...]}{(2j)!}$ (82)
[…]
so that the final expression for the world function can be written concisely as follows,
$\Omega(t,T;\ell)=\Omega(0)+\cosh(\ell D_{\alpha}[G^{-1}\Omega(\alpha)])+\ell\sinh(\ell D_{\alpha}[G^{-1}(G^{-1})^{\prime}\Omega(\alpha)])$ (84)
Acknowledgements
Mayank thanks the Indian Institute of Technology (IIT) Madras and the Ministry of Human Resources and Development (MHRD), India for financial support.